# Causally Reliable Concept Bottleneck Models

### v2

@@ A.1 Pearl’s framework of causality
Contemporary research in causal inference and causal machine learning predominantly builds on the framework introduced by Pearl (2009) (see also Pearl (2019)). This framework centers on an agent’s ability to reason about underlying causal mechanisms, going beyond mere statistical associations observed in data. Pearl formalizes this capacity through the notion of answering different types of what-if questions, structured into a three-level hierarchy (see Figure 6).
Figure 6: Pearl’s Hierarchy as represented in Pearl & Mackenzie (2018).
  • Level 1: Observational questions (“what is the value of $Y$ if I observe $X=x$?”). These questions rely purely on statistical associations within the data. They can be answered using standard tools such as conditional probabilities or expectations, with no need for causal assumptions.
  • Level 2: Interventional questions (“what is the value of $Y$ if I do $X=x$?”). These questions focus on the effects of actively intervening on variables, rather than passively observing them. The expression $\text{do}(X=x)$ denotes such an intervention, where $X$ is externally set to $x$, overriding its natural causes.
@@
  • Level 3: Counterfactual questions (“What would have been the value of $Y$ if I had observed $X=x^{\prime}$ instead of $X=x$?”). These questions focus on states of affairs alternative to the actual reality. They constitute the upper layer of Pearl’s hierarchy and their computation requires a detailed knowledge of the causal mechanisms relating the variables of the target-problem.
While standard machine learning models  [^3] This excludes models specifically designed for causal inference, such as those in the field of causal machine learning (Kaddour et al., 2022). can generally handle only observational questions, reasoning at interventional and counterfactual levels necessitates dedicated causal models and inference methods (Pearl, 2009; Peters et al., 2017).
@@ A.2 Causal opacity
Note that causal transparency does not presuppose or imply causal reliability, the two issues, although related, are indeed very different. Consider for example a concept-based model $\mathcal{M}_{1}$ represented in the figure below. This model is trained to predict whether a given colored image represents an even or odd number (variable $P$, for “parity”) passing through a bottleneck of two interpretable concepts, namely “number” ($N$) and “color” ($C$). The structure of $\mathcal{M}_{1}$’s decision-making process is causally-transparent, including two edges connecting the final task $P$ with $N$ and $C$ respectively. However, this causal structure is not consistent with the causal structure of the world, for which $C$ and $P$ are clearly independent concepts Therefore, $\mathcal{M}_{1}$ cannot be considered causally reliable, although it is causally transparent.
(a) Causal transparent graph $N$ $C$ $P$ (b) Causal reliable graph $N$ $C$ $P$
@@ Appendix B Concept discovery details
CBMs-like architectures tipically rely on labeled data for each concept, a costly process that can hinder their practical adoption. Label-free Concept Bottleneck Models (Label-free CBMs) (Oikarinen et al., 2023) address this issue by automatically generating concepts and assign concept labels using pre-trained models.
In our paper, we apply an approach similar to the one followed in (Oikarinen et al., 2023) to the Siim-Pneumothorax dataset, which lacks concepts and concept annotations. Using GPT-4o, we first generate candidate concepts with a specific prompt (see App. C.3), and apply a multi-stage filtering process:
@@
  • filter out concepts too similar to class labels or to each other. Specifically, we use the CXR-CLIP model—a CLIP-based model pretrained on medical imaging datasets(Johnson et al., 2019; Irvin et al., 2019; Wang et al., 2017; You et al., 2023)—to encode both concepts and class labels, and discard any concept with cosine similarity > 0.9 to either a class label or another concept;
@@ C.1 Causal discovery algorithms
Algorithms for addressing the causal discovery problem can be broadly classified into two main categories (Peters et al., 2017): [^4] Here, we restrict our discussion to causal discovery algorithms that assume the set of observed variables sufficiently captures the relevant causal influences. For more complex scenarios, see, e.g., (Peters et al., 2017).
@@
  • Score-based methods. These methods define a scoring function over potential graph structures and search for the graph that maximizes it, often using criteria such as the Bayesian Information Criterion (BIC). The results from score-based methods are often comparable to those from independence-based approaches, as graphs that violate conditional independencies tend to result in poor model fits.
For our experiments, we adopted the Greedy Equivalence Search (GES) algorithm, a score-based method that performed well in our setting, as shown in Table 8 in App. G.4.
GES is a two-phase greedy algorithm that searches over equivalence classes of DAGs (CPDAGs). It begins with the empty graph and iteratively adds edges that yield the greatest improvement in the scoring function. Once a local optimum is reached—where no addition improves the score—the algorithm enters a backward phase, greedily removing edges that most increase the score (Chickering, 2002). The core idea is to navigate the space of CPDAGs through local transformations (edge additions and deletions), using the score as a guide to optimize structure learning.
In our implementation, we use the GES algorithm provided by the causallearn Python library, employing the Bayesian Dirichlet equivalent uniform (BDeu) scoring criterion for discrete variables (Heckerman et al., 1995; Chickering, 2002).
@@ C.2 LLM and RAG
LLMs (Brown et al., 2020; Jiang et al., 2024) are capable of answering complex queries without additional training. However, they can be unreliable and prone to hallucinations(Ji et al., 2023; Zhang et al., 2023). Retrieval Augmented Generation (RAG) (Lewis et al., 2020) addresses this by retrieving relevant textual information and appending it to the query, thereby improving the accuracy and reliability of LLM outputs. In the context of structural learning, LLMs have been utilized to construct causal graphs by employing ad-hoc prompts specifically designed to condition the model for answering causal queries (Antonucci et al., 2023; Long et al., 2023; Zhang et al., 2024). To enhance the causal graph discovery process, we utilize an LLM integrated with a RAG to either direct or eliminate edges that remain undirected by the causal discovery algorithm. The information retrieval process is outlined as follows:
@@
    Document Retrieval: Given a causal query (e.g., “Is lung cancer influenced by smoking?”), we first retrieve a set of documents from the web. In our experiments, we employed both the DuckDuckGo search engine for web pages and Arxiv for relevant paper abstracts. From each source, we retrieve the top 10 documents. For certain datasets (ColorMNIST, CelebA, and Sachs), local documents were used due to the unavailability or inaccessibility of information online (e.g., the Sachs paper). Each retrieved document is then segmented into smaller pieces, referred to as chunks, using a sliding window approach that samples a 512-token chunk every 128 tokens.
@@
    Ranking: The causal query is transformed by the LLM using a query transformation approach (Gao et al., 2023), which improves the semantic alignment of the query with the relevant document chunks. After transformation, all the retrieved chunks and the modified causal query are processed by a sentence transformer (Reimers, 2019), specifically the multi-qa-mpnet-base-dot-v1 model. At this stage, cosine similarity is computed between the embedded transformed causal query and each embedded chunk.
@@ Appendix D C2BMs detailed architecture
In this appendix, we provide a detailed description of the proposed C2BM model training and functioning, using the Asia dataset as an illustrative example (Fig. 7). We assume the following information is available:
@@
  • A training dataset $D=\{\textbf{x}_{i},\textbf{v}_{i}\}_{i=1}^{n}$, where each sample is annotated with the values of all endogenous variables, including the target variable Dyspnea and all other relevant concepts.
@@
These elements can be either provided by human experts or generated using the pipeline we propose in the paper, as described in Fig. 2. Consider, for example, Dyspnea as our binary target variable serving as the task. The proposed approach follows four main steps (Fig. 7), as detailed below:
@@
    Sub-causal Graph Selection. We extract from the DAG $\mathbf{G}$ only the variables that are ancestors of the task, i.e., there exists a causal path in $\mathbf{G}$ connecting the variable node to the task. In the case of the Asia dataset, all the variables are connected to the task, thus the graph is left unaltered.
@@
    Exogenous embeddings. Each concept (including the task) is assumed to have an associated latent factor, which is represented as an embedding (the grey encoder symbols in Fig. 7) learned from the input using a dedicated neural encoder. In our implementation, these are implemented as MLPs, preceded by a dataset-specific feature extractor, e.g., a CNN for image data (as detailed in the App. E).
@@
    Structural Equation Modeling. We model the structural relationships between parent nodes and their child nodes using linear equations. The weights of these equations are predicted by MLPs, which take as input the embedding of the child node produced in the previous step. Applying these functions we can derive the normalized logits of a child node from the ones of its parents. For the nodes that lie in the roots of the causal graph (Been in Asia, Smoker), their logits are obtained directly from the corresponding exogenous embeddings.
@@ D.1 C2BM as an universal approximator
We show that C2BM can predict any endogenous variable as any DNN. Assume the exogenous encoder $\mathbf{g}$ is a DNN (or any universal approximator) and that endogenous variables are represented as logits (the extension to fuzzy or Boolean values is trivial). We consider four cases:
@@
Hence, by recursively composing universal approximators along the causal graph, C2BM can approximate any mapping from the input to endogenous variables. This holds for any graph structure, establishing C2BM as a universal approximator. ∎
@@ D.2 C2BM interpretability
Figure 9: Visualization example of information propagation (Asia dataset). The parent weight (predicted by the meta-model) is visualized next to each edge.
@@ Appendix E Dataset details
The MNIST dataset (LeCun et al., 2010) is a large collection of freely available grayscale images of handwritten digits. It consists of $60000$ training images and $10000$ test images, both drown from the same distribution. Each image is labeled with the digit it represents. For our experiments, we download the original train and test dataset using the torchvision library (maintainers & contributors, 2016) and reserve $10\%$ of the training set for validation. Additionally, we colorize each image based on its digit. Specifically, in the in-distribution setting (experiment in Sec.5.1), color is randomly assigned to each image in the training, validation, and test sets with equal probability over red, green, and blue. In the out-of-distribution setting (experiment in Sec.5.3), images in the training and validation sets with odd digits are colored green, while the remaining images are colored red or blue with equal probability. In the test set, images with even digits are colored green, and the remaining ones are colored red or blue with equal probability. For both the versions of this dataset, the following concepts are considered: Number, Color, and Parity (task). Finally, the images are preprocessed using a pre-trained ResNet-18 model with default weights from the torchvision library.
@@ E.2 Bayesian networks
For our experiments, we use synthetic datasets sampled from discrete Bayesian networks available in the bnlearn repository https://www.bnlearn.com/bnrepository/. A Bayesian network is a probabilistic graphical model consisting of a DAG, where nodes correspond to random variables, and each node is associated with a conditional probability distribution (CPD). This CPD defines the probability of the node’s value, given the values of its parent nodes in the network (Sharma et al., 2020). From the bnlearn repository, we select Bayesian networks with different dimensions and domains: Asia (Lauritzen & Spiegelhalter, 1988), a small network focused on lung disease with $8$ nodes and $8$ edges; Sachs (Sachs et al., 2005), a widely-used network modeling the relationships between protein and phospholipid expression levels in human cells with $11$ nodes and $17$ edges; Insurance (Binder et al., 1997), a network for evaluating car insurance risks with $27$ nodes and $52$ edges; Alarm (Beinlich et al., 1989), a network designed to provide an alarm message system for patient monitoring with $37$ nodes and $46$ edges; Hailfinder (Abramson et al., 1996), a network designed to forecast severe summer hail in northeastern Colorado with $56$ nodes and $66$ edges. For each network, we generate $10000$ samples and create training, validation, and test datasets using a $70\%-10\%-20\%$ split.
While node values can be used as concept annotations ($\mathbf{v}$), input features ($\mathbf{x}$) are absent. To generate them and make the datasets applicable to concept-based architectures, we flatten the concept values and process them with a simple autoencoder (MSE loss) comprising 2 encoder layers and 2 decoder layers (the latent dimension is adjusted based on the number of nodes: Asia-32, Sachs-32, Insurance-32, Alarm-64, Hailfinder-128. Embeddings are further transformed so that each sample is a mixture composed of $50\%$ original data and $50\%$ noise, with the noise drawn from a standard normal distribution. Finally, the output is standardized. The goal of these transformations is to make the inputs non-trivial representations of the concepts (the network nodes), forcing architectures to learn how to identify and retrieve the underlying concepts from the preprocessed data.
@@ E.3 CelebA
CelebA (Liu et al., 2015) is a large-scale face attributes dataset with more than $200.000$ celebrity images divided into training, validation and test set with $40$ binary attribute annotations. For our experiments, we first downloaded all the splits from the project website https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html.
@@ E.4 Siim-pneumothorax
This dataset is derived from the publicly available chest radiograph dataset provided by the National Institutes of Health (NIH) and it contains chest x-ray images with binary annotations indicating the presence or the absence of Pneumothorax. For our experiments, we use in particular the training annotations available on Kaggle (https://www.kaggle.com/competitions/siim-acr-pneumothorax-segmentation) and the corresponding training images from https://www.kaggle.com/datasets/abhishek/siim-png-images. The dataset is then split into training, validation, and test sets using a traditional $70\%-10\%-20\%$ partition. As it does not include concept annotations, we followed a procedure inspired by the methodology in Oikarinen et al. (2023) to generate concepts and their corresponding annotations. More details can be found in App. B.
E.5 CUB
This dataset is derived from the publicly available Caltech-UCSD Birds-200-2011 (CUB) (He & Peng, 2019) dataset, which is widely used in the CBM community (Koh et al., 2020; Zarlenga et al., 2022). It contains 11,788 images across 200 bird categories, with 5,994 images for training and 5,794 images for testing. Each image is annotated in detail, including 312 binary attributes. For our experiments, we downloaded the dataset from https://data.caltech.edu/records/65de6-vp158 and further split the training set such that $10\%$ is used for validation. We then selected the 112 most frequently activated binary attributes to serve as our concepts of interest, as considered in Zarlenga et al. (2022).

### v3

@@ A.1 Pearl’s framework of causality
Contemporary research in causal inference and causal machine learning predominantly builds on the framework introduced by Pearl2009 (see also Pearl2019). This framework centers on an agent’s ability to reason about underlying causal mechanisms, going beyond mere statistical associations observed in data. Pearl formalizes this capacity through the notion of answering different types of what-if questions, structured into a three-level hierarchy (see Figure 6).
Figure 6: Pearl’s Hierarchy as represented in Pearl2018.
  • Level 1: Observational questions (“what is the value of $Y$ if I observe X=xX=x?”). These questions rely purely on statistical associations within the data. They can be answered using standard tools such as conditional probabilities or expectations, with no need for causal assumptions.
  • Level 2: Interventional questions (“what is the value of $Y$ if I do X=xX=x?”). These questions focus on the effects of actively intervening on variables, rather than passively observing them. The expression $\text{do}(X=x)$ denotes such an intervention, where $X$ is externally set to $x$, overriding its natural causes.
@@
  • Level 3: Counterfactual questions (“What would have been the value of $Y$ if I had observed $X=x^{\prime}$ instead of X=xX=x?”). These questions focus on states of affairs alternative to the actual reality. They constitute the upper layer of Pearl’s hierarchy and their computation requires a detailed knowledge of the causal mechanisms relating the variables of the target-problem.
While standard machine learning models  [^5] This excludes models specifically designed for causal inference, such as those in the field of causal machine learning (kaddour2022causal). can generally handle only observational questions, reasoning at interventional and counterfactual levels necessitates dedicated causal models and inference methods (Pearl2009; peters2017elements).
@@ A.2 Causal opacity
Note that causal transparency does not presuppose or imply causal reliability, the two issues, although related, are indeed very different. Consider for example a concept-based model $\mathcal{M}_{1}$, where the graph representing how the concepts are connected to the final task is shown in the left panel of the figure below (a). This model is trained to predict whether a given colored image represents an even or odd number (variable $P$, for “parity”) passing through a bottleneck of two interpretable concepts, namely “number” ($N$) and “color” ($C$). The structure of $\mathcal{M}_{1}$’s decision-making process is causally-transparent, including two edges connecting the final task $P$ with $N$ and $C$ respectively. However, this causal structure is not consistent with the causal structure of the world (b), for which $C$ and $P$ are clearly independent concepts Therefore, $\mathcal{M}_{1}$ cannot be considered causally reliable, although it is causally transparent.
(a) Causal transparent graph NN $C$ $P$ (b) Causal reliable graph NN $C$ $P$
@@ Appendix B Concept discovery details
CBMs-like architectures tipically rely on labeled data for each concept, a costly process that can hinder their practical adoption. Label-free Concept Bottleneck Models (Label-free CBMs) (oikarinen2023) address this issue by automatically generating concepts and assign concept labels using pre-trained models.
In our paper, we apply an approach similar to the one followed in oikarinen2023 to the Siim-Pneumothorax dataset, which lacks concepts and concept annotations. Using GPT-4o, we first generate candidate concepts with a specific prompt (see App. C.3), and apply a multi-stage filtering process:
@@
  • filter out concepts too similar to class labels or to each other. Specifically, we use the CXR-CLIP model – a CLIP-based model pretrained on medical imaging datasets (johnson2019; irvin2019; wang2017; you2023) – to encode both concepts and class labels, and discard any concept with cosine similarity > 0.9 to either a class label or another concept;
@@ C.1 Causal discovery algorithms
Algorithms for addressing the causal discovery problem can be broadly classified into two main categories (peters2017elements): [^6] Here, we restrict our discussion to causal discovery algorithms that assume the set of observed variables sufficiently captures the relevant causal influences. For more complex scenarios, see, e.g., (peters2017elements).
@@
  • Score-based methods. These methods define a scoring function over potential graph structures and search for the graph that maximizes it, often using criteria such as the Bayesian Information Criterion (BIC) (peters2017elements). The results from score-based methods are often comparable to those from independence-based approaches, as graphs that violate conditional independencies tend to result in poor model fits.
For our experiments, we adopted the Greedy Equivalence Search (GES) algorithm (chickering2002), a score-based method that performed well in our setting, as shown in Table 8 in App. G.3.
GES is a two-phase greedy algorithm that searches over equivalence classes of DAGs (CPDAGs). It begins with the empty graph and iteratively adds edges that yield the greatest improvement in the scoring function. Once a local optimum is reached — where no addition improves the score — the algorithm enters a backward phase, greedily removing edges that most increase the score. The core idea is to navigate the space of CPDAGs through local transformations (edge additions and deletions), using the score as a guide to optimize structure learning.
In our implementation, we use the GES algorithm provided by the causal-learn Python library (zheng2024), employing the Bayesian Dirichlet equivalent uniform (BDeu) scoring criterion for discrete variables (heckerman1995learning; chickering2002).
@@ C.2 LLM and RAG
LLMs (brown2020language; jiang2024mixtral) are capable of answering complex queries without additional training. However, they can be unreliable and prone to hallucinations (ji2023survey; zhang2023siren). Retrieval Augmented Generation (RAG) (lewis2020retrieval) addresses this by retrieving relevant textual information and appending it to the query, thereby improving the accuracy and reliability of LLM outputs. In the context of structural learning, LLMs have been utilized to construct causal graphs by employing ad-hoc prompts specifically designed to condition the model for answering causal queries (antonucci2023zero; long2023can; zhang2024causal). To enhance the causal graph discovery process, we utilize an LLM integrated with a RAG to either direct or eliminate edges that remain undirected by the causal discovery algorithm. The information retrieval process is outlined as follows:
@@
    Document Retrieval: Given a causal query (e.g., “Is lung cancer influenced by smoking?”), we first retrieve a set of documents from the web. In our experiments, we employed both the DuckDuckGo search engine for web pages and Arxiv for relevant paper abstracts. From each source, we retrieve the top 10 documents. For certain datasets (cMNIST, CelebA, and Sachs), local documents were used due to the unavailability or inaccessibility of information online (e.g., the Sachs paper). Each retrieved document is then segmented into smaller pieces, referred to as chunks, using a sliding window approach that samples a 512-token chunk every 128 tokens.
@@
    Ranking: The causal query is transformed by the LLM using a query transformation approach (gao2023retrieval), which improves the semantic alignment of the query with the relevant document chunks. After transformation, all the retrieved chunks and the modified causal query are processed by a sentence transformer (reimers2019sentence), specifically the multi-qa-mpnet-base-dot-v1 model. At this stage, cosine similarity is computed between the embedded transformed causal query and each embedded chunk.
@@ Appendix D C2BMs detailed architecture
In this appendix, we provide a detailed description of the proposed C2BM model training and functioning, using the Asia dataset as an illustrative example and Dyspnea as the task (Fig. 7). We assume the following information is available:
@@
  • A training dataset $D=\{\textbf{x}_{i},\textbf{v}_{i}\}_{i=1}^{n}$, where each sample is annotated with the values of all endogenous variables, i.e., the target variable Dyspnea and all preceding binary concepts.
@@
These elements can be either provided by human experts or generated using the pipeline we propose in the paper, as described in Fig. 2. The proposed approach follows four main steps (Fig. 7), as detailed below:
@@
    Sub-causal Graph Selection. We extract from the DAG $\mathbf{G}$ only the variables that are ancestors of the task, that is, variables for which there exists a causal path in $\mathbf{G}$ connecting the variable node to the task. In the case of the Asia dataset, the variable X-ray anomalies is discarded because it is not an ancestor of the task.
@@
    Exogenous embeddings. Each variable (including the task) is assumed to have an associated latent factor, which is represented as an embedding (the grey encoder symbols in Fig. 7) learned from the input using a dedicated neural encoder. In our implementation, these are implemented as MLPs, preceded by a dataset-specific feature extractor, e.g., a CNN for image data (as detailed in the App. E).
@@
    Structural Equation Modeling. We model the structural relationships between parent nodes and their child nodes using linear equations. The weights of these equations are predicted by separate hypernetworks, implemented as MLPs, which take as input the embedding of the child node produced in the previous step. Applying these functions we can derive the normalized logits of a child node from the ones of its parents. For the nodes that lie in the roots of the causal graph (Been in Asia, Smoker), their logits are obtained directly from the corresponding exogenous embeddings.
@@ D.1 C2BM as an universal approximator
We show that C2BM can predict any endogenous variable as any DNN. Assume the exogenous encoder $\mathbf{g}(\cdot)$ is a DNN (or any universal approximator) and that endogenous variables are represented as logits (the extension to fuzzy or Boolean values is trivial). We consider four cases:
@@
Hence, by recursively composing universal approximators along the causal graph, C2BM can approximate any mapping from the input to endogenous variables. This holds for any graph structure, establishing C2BM as a universal approximator.
@@ D.2 C2BM interpretability
Figure 9: Visualization example of information propagation (Asia dataset). The parent weight (predicted by the hypernetwork) is visualized next to each edge.
@@ Appendix E Dataset details
The MNIST dataset (lecun2010) is a large collection of freely available grayscale images of handwritten digits. It consists of $60000$ training images and $10000$ test images, both drown from the same distribution. Each image is labeled with the digit it represents. For our experiments, we download the original train and test dataset using the torchvision library (marcel2010) and reserve $10\%$ of the training set for validation. Additionally, we colorize each image based on its digit. Specifically, in the in-distribution setting (experiment in Sec.5.1), color is randomly assigned to each image in the training, validation, and test sets with equal probability over red, green, and blue. In the out-of-distribution setting (experiment in Sec.5.3), images in the training and validation sets with odd digits are colored green, while the remaining images are colored red or blue with equal probability. In the test set, images with even digits are colored green, and the remaining ones are colored red or blue with equal probability. For both the versions of this dataset, the following concepts are considered: Number, Color, and Parity (task). Finally, the images are preprocessed using a pre-trained ResNet-18 model with default weights from the torchvision library.
@@ E.2 Bayesian networks
For our experiments, we use synthetic datasets sampled from discrete Bayesian networks available in the bnlearn repository (https://www.bnlearn.com/bnrepository/). A Bayesian network is a probabilistic graphical model consisting of a DAG, where nodes correspond to random variables, and each node is associated with a conditional probability distribution (CPD). This CPD defines the probability of the node’s value, given the values of its parent nodes in the network (sharma2020). From the bnlearn repository, we select Bayesian networks with different dimensions and domains: Asia (asia), a small network focused on lung disease with $8$ nodes and $8$ edges; Sachs (sachs2005causal), a widely-used network modeling the relationships between protein and phospholipid expression levels in human cells with $11$ nodes and $17$ edges; Insurance (binder1997), a network for evaluating car insurance risks with $27$ nodes and $52$ edges; Alarm (alarm), a network designed to provide an alarm message system for patient monitoring with $37$ nodes and $46$ edges; Hailfinder (abramson1996), a network designed to forecast severe summer hail in northeastern Colorado with $56$ nodes and $66$ edges. For each network, we generate $10000$ samples and create training, validation, and test datasets using a $70\%-10\%-20\%$ split.
While node values can be used as concept annotations ($\mathbf{v}$), input features ($\mathbf{x}$) are absent. To generate them and make the datasets applicable to concept-based architectures, we flatten the concept values and process them with a simple autoencoder (MSE loss) comprising 2 encoder layers and 2 decoder layers (the latent dimension is adjusted based on the number of nodes: Asia-32, Sachs-32, Insurance-32, Alarm-64, Hailfinder-128). Embeddings are further transformed so that each sample is a mixture composed of $50\%$ original data and $50\%$ noise, with the noise drawn from a standard normal distribution. Finally, the output is standardized. The goal of these transformations is to make the inputs non-trivial representations of the concepts (the network nodes), forcing architectures to learn how to identify and retrieve the underlying concepts from the preprocessed data.
@@ E.3 CelebA
CelebA (liu2015) is a large-scale face attributes dataset with more than $200.000$ celebrity images divided into training, validation and test set with $40$ binary attribute annotations. For our experiments, we first downloaded all the splits from the project website https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html.
@@ E.4 Siim-pneumothorax
This dataset is derived from the publicly available chest radiograph dataset provided by the National Institutes of Health (NIH) and it contains chest x-ray images with binary annotations indicating the presence or the absence of Pneumothorax. For our experiments, we use in particular the training annotations available on Kaggle (https://www.kaggle.com/competitions/siim-acr-pneumothorax-segmentation) and the corresponding training images from https://www.kaggle.com/datasets/abhishek/siim-png-images. The dataset is then split into training, validation, and test sets using a traditional $70\%-10\%-20\%$ partition. As it does not include concept annotations, we followed a procedure inspired by the methodology in oikarinen2023 to generate concepts and their corresponding annotations. More details can be found in App. B.
E.5 CUBC
This dataset is derived from the publicly available Caltech-UCSD Birds-200-2011 (CUB) (he2019fine) dataset, which is widely used in the CBM community (koh2020; zarlenga2022concept). It contains 11,788 images across 200 bird categories, with 5,994 images for training and 5,794 images for testing. Each image is annotated in detail, including 312 binary attributes. For our experiments, we downloaded the dataset from https://data.caltech.edu/records/65de6-vp158 and further split the training set such that $10\%$ is used for validation. We then selected the 112 most frequently activated binary attributes to serve as our concepts of interest, as considered in zarlenga2022concept.